# Revisiting Data Challenges of Computational Pathology: A Pack-based Multiple Instance Learning Framework

### v1

@@ Abstract
Computational pathology (CPath) digitizes pathology slides into whole slide images (WSIs), enabling analysis for critical healthcare tasks such as cancer diagnosis and prognosis. However, WSIs possess extremely long sequence lengths (up to 200K), significant length variations (from 200 to 200K), and limited supervision. These extreme variations in sequence length lead to high data heterogeneity and redundancy. Conventional methods often compromise on training efficiency and optimization to preserve such heterogeneity under limited supervision. To comprehensively address these challenges, we propose a pack-based MIL framework. It packs multiple sampled, variable-length feature sequences into fixed-length ones, enabling batched training while preserving data heterogeneity. Moreover, we introduce a residual branch that composes discarded features from multiple slides into a hyperslide which is trained with tailored labels. It offers multi-slide supervision while mitigating feature loss from sampling. Meanwhile, an attention-driven downsampler is introduced to compress features in both branches to reduce redundancy. By alleviating these challenges, our approach achieves an accuracy improvement of up to 8% while using only 12% of the training time in the PANDA(UNI). Extensive experiments demonstrate that focusing data challenges in CPath holds significant potential in the era of foundation models. The code is here.
@@ 1 Introduction
Computational pathology (CPath) Song et al. (2023), Cifci et al. (2023) represents a rapidly evolving interdisciplinary research domain that integrates advanced computer vision techniques and pathology to facilitate accurate and efficient interpretation of histopathological images. Central to CPath are whole slide images (WSIs, slides), digitized pathology slides with gigapixel resolution often exceeding billions of pixels. It enables comprehensive microscopic analysis to support critical healthcare tasks such as cancer sub-typing Ilse et al. (2018), Zhang et al. (2022), Tu et al. (2022), grading Bulten et al. (2022), and prognosis Wen et al. (2023), Yao et al. (2020). Patching strategies help researchers effectively process these gigapixel images within hardware constraints. However, patches derived from WSIs present following challenges: as shown in Fig.1, 1) they possess extremely long sequence lengths (up to 200K) and significant sequence length variations (from 200 to 200K). Such extreme distributions in sequence length, coupled with diverse morphological characteristics, contribute to data heterogeneity, which is substantial for CPath tasks. 2) and introduce input redundancy challenges for CPath algorithms. 3) Moreover, due to the gigapixel resolution and specialization, WSIs typically have only slide-level annotations, lacking more supervision that matches the complex input.
@@
Current two-stage multiple instance learning (MIL) Maron & Lozano-Pérez (1997) paradigm Lu et al. (2021) is a compromise resulting from high data heterogeneity and limited supervision. This paradigm employs a pre-trained encoder to extract offline patch (instance) features, and uses a MIL model to produce slide-level (bag-level) results. Due to data challenges, it suffers from training inefficiency and instability. Specifically, with significant variations in sequence length across slides, mainstream methods typically process data with a batchsize of 1 during training Shao et al. (2021), Li et al. (2024c). While these approaches preserve whole-slide heterogeneity, training with a batchsize of 1 is inefficient (e.g., training TransMIL Shao et al. (2021) on the PANDA Bulten et al. (2022) dataset requires over 50 RTX3090 GPU-hours) and may yield suboptimal performance Koga et al. (2025). A few methods Campanella et al. (2019), Liu et al. (2024b) attempt to enable batched training by sampling or padding all sequences to a uniform length; however, this approach can lead to a loss of data heterogeneity and important features, especially affecting complex methods and tasks.
@@
To comprehensively address three data challenges, we propose a novel pack-based MIL framework. Inspired by recent advancements in large model Pouransari et al. (2024), Krell et al. (2021), Dehghani et al. (2023), Wang et al. (2024a), it packs multiple variable-length sequences into a single fixed-length sequence to enable batched training while preserving data heterogeneity. However, leveraging packing strategies for effective batched training in CPath is far from straightforward. The excessive length of packed sequences hinders training, necessitating patch sampling, which still leads to feature loss. Therefore, we split the input features into main and residual branches, packing the retained and discarded features, respectively, to minimize sampling-induced feature loss. In the main branch, we employ masks to maintain the independence of different slides within a pack. Conversely, the residual branch treats discarded features from multiple slides in the same pack as a single hyperslide. To train this hyperslide effectively, we introduce task-specific hyperslide labels and loss functions. Crucially, this approach effectively offers multi-slide supervision.
@@
While some outstanding works have explored supplementary supervision Zhang et al. (2022), Shao et al. (2023), Brussee et al. (2025), Fang et al. (2024), most focus on intra-slide modeling (e.g., instance-level or pseudo-bag), neglecting inter-slide relationships. Pathology slides from the same spatial and tissue origin exhibit consistent morphological characteristics Lin et al. (2025), Chen et al. (2022a), Kaczmarzyk et al. (2024). Learning inter-slide correlations allows the hyperslide to provide the model with a more comprehensive perspective, enabling the discovery of more generalizable pathological features. Furthermore, we propose an attention-driven downsampler to compress features for reducing input redundancy within both branches. To validate our framework, we conducted extensive experiments using features from foundation models. Results demonstrate that our approach consistently improves multiple baselines by effectively mitigating the data challenges inherent in CPath. Specifically, it delivers substantial performance gains (e.g., +11% accuracy on PANDA) while improving training efficiency ($\sim 8\times$ speedup on PANDA). Our contributions are:
@@ 2 Related Works
Supervision in Computational Pathology. Recent CPath advancements leverage MIL to reduce annotation burden. Using only slide labels, MIL has evolved with mechanisms like attention Ilse et al. (2018), Tang et al. (2023), clustering Lin et al. (2023), Transformers Shao et al. (2021), and GNNs Wang et al. (2021), Eastwood et al. (2023) to improve interpretability, localization, and accuracy. Complementing pure MIL methods, pseudo-labeling strategies have emerged as powerful auxiliary techniques, encompassing instance-level pseudo-labeling Qu et al. (2022a), knowledge distillation frameworks Zhang et al. (2022), Qu et al. (2022b), limited pathologist patch annotations Koga et al. (2025), weak regional annotations Wang et al. (2022), and semi-supervised consistency regularization Jiang et al. (2023). These hybrid strategies effectively generate additional supervision to refine instance predictions and leverage unlabeled data, boosting performance and data efficiency. While some studies Liu et al. (2024a), Ouyang et al. (2024) explore mixup-like data augmentation between WSI pairs, supervision leveraging relationships across multiple slides is still largely unexplored.
Batchsize in Computational Pathology. Batchsize is a crucial hyperparameter in deep learning. However, its exploration in CPath remains limited, primarily due to the computational demands of gigapixel WSIs and the inherent data heterogeneity within each slide. Consequently, mainstream slide-level MIL methods typically adopt a batchsize of 1 Jaume et al. (2024), Li et al. (2024a), Shi et al. (2024), Li et al. (2024c), Song et al. (2024), Zhang et al. (2024b), Li et al. (2024b), Fourkioti et al. (2023), Tang et al. (2025b; a). For instance, RRTMIL Tang et al. (2024) utilizes slide-wise regional and cross-region self-attention to capture patch ordinality and heterogeneity within each slide, necessitating a batchsize of 1 to maintain intra-slide relationships. Despite its prevalence, this practice often results in training instability and slow convergence, prompting methods such as gradient accumulation over multiple slides Koga et al. (2025), Zhang et al. (2025) or instance-level sampling strategies that select fixed-size subsets of patches per slide to mitigate computational overhead and improve learning stability Campanella et al. (2019), Liu et al. (2024b). Current slide-level methods treat batched inputs as an efficiency trade-off rather than a genuinely effective training strategy. In this paper, we investigate the benefits of batched training for slide‑level prediction and explore practical guidelines for its implementation in CPath. Appendix E gives more discussion.
@@ 3 Method
3.1 Preliminary: MIL-based Computational Pathology
Figure 2: Impact of data heterogeneity on CPath.
@@ 3.1 Preliminary: MIL-based Computational Pathology
Histopathological WSIs are often gigapixel resolution, making direct processing impractical. Current approaches typically use weakly supervised MIL, where a WSI is treated as a bag $\mathcal{B}=\{x_{1},\dots,x_{N}\}$ of instances. During training, only a slide‑level label $y$ is available. Each instance $x_{i}$ is encoded to an embedding $h_{i}=f(x_{i})$ using a offline feature extractor $f$. An aggregation function $\Gamma(\cdot)$ combines instance embeddings $\{h_{i}\}_{i=1}^{N}$ into a bag-level representation $z=\Gamma_{\theta}(\sum_{i=1}^{N}h_{i})$. This representation $z$ is then used by a classifier $g_{\phi}$ to predict the slide label $p(y\,|\,\mathcal{B})=g_{\phi}(z)$. The significant variation in instance count $N$ across WSIs contribute to data heterogeneity. This heterogeneity along with its associated spatial and morphological context, is crucial for CPath, as shown in Fig.2. Specifically, when all instances are randomly sampled to a fixed-length sequence, a latest method like RRTMIL Tang et al. (2024) exhibits consistent performance degradation on multiple benchmarks, particularly for complex tasks like survival analysis. To maintain data heterogeneity, current methods necessitates training with a batchsize of 1, resulting in noisy gradient estimates and optimization instability.
Figure 3: Left: Overview of the proposed pack-based MIL training framework. Right: llustration of Attention-driven Downsampler (ADS).
3.2 Pack-based MIL Framework
@@
CPath poses significant data challenges, including high data heterogeneity, redundancy and limited supervision, which hinder mainstream MIL method. To overcome these issues, we propose a pack-based MIL framework, named PackMIL. As illustrated in Fig.3, the proposed method effectively employs a packing operation to maintain data heterogeneity and allow batched input, and it further provides multi-slide supervision, leading to more effective and stable training. Consider a mini‑batch of $B$ bags, the bb‑th bag provides a set of instance embeddings $\mathcal{B}_{b}=\{h_{bi}\}_{i=1}^{N_{b}}$ where $h_{bi}\!\in\!\mathbb{R}^{D}$. However, the feature sequences extracted from gigapixel WSIs can be prohibitively long, making their direct inclusion into fixed-length packs impractical due to memory and computational limitations during training. To address this, we employ stochastic instance-level sampling with ratio $r\in(0,1)$. The subsets $\mathcal{R}_{b}$ and $\mathcal{D}_{b}$ are formed as:
$m_{bi}\sim\operatorname{Bernoulli}(1-r),\quad{\mathcal{R}}_{b}=\{\,h_{bi}\mid m_{bi}=1\,\},\quad\mathcal{D}_{b}=\{\,h_{bi}\mid m_{bi}=0\,\},$ (1)
@@
We introduce the Residual Branch to mitigate feature loss during sampling and establish hyperslide supervision. The retained set $\mathcal{R}$ is forwarded in main branch, while discarded set $\mathcal{D}$ is forwarded in residual branch. For each bag $b$, the corresponding sets $\mathcal{R}_{b}$ and $\mathcal{D}_{b}$ are processed to produce downsampled feature sets, denoted as $\tilde{\mathcal{R}}_{b}$ and $\tilde{\mathcal{D}}_{b}$, respectively. To maintain data heterogeneity for batch training, instances from sampled sets of each bag are sequentially arranged into fixed-length packs of size $L$. This packing operation $\operatorname{PACK}(\cdot)$ processes all bags in the mini-batch to form:
@@
$P^{\mathrm{main}}=\operatorname{PACK}\Bigl(\bigcup_{b=1}^{B}\tilde{\mathcal{R}}_{b},\,L\Bigr)\in\mathbb{R}^{B^{\prime}\times L\times D},P^{\mathrm{res}}=\operatorname{PACK}\Bigl(\bigcup_{b=1}^{B}\tilde{\mathcal{D}}_{b},\,L\Bigr)\in\mathbb{R}^{B^{\prime\prime}\times L\times D},$ (2)
@@
We also incorporate an attention-driven downsampler (ADS) module for handling input redundancy between the two branches. This module fuses features from instances in $\mathcal{R}_{b}$ and $\mathcal{D}_{b}$ to generate more compact and informative feature representations. Each set $\mathcal{S}\in\{\mathcal{R}_{b},\mathcal{D}_{b}\}$ is downsampled by a factor $k\in\mathbb{N}$ using an ADS module, yielding $\tilde{\mathcal{S}}=\operatorname{ADS}(\mathcal{S};k)$ with size $|\tilde{\mathcal{S}}|=\lceil|\mathcal{S}|/k\rceil$.
@@
After packing, the main and residual branches are processed independently while sharing the same MIL model weight. In the main branch, each pack contains instances originating from $B$ distinct bags. We use a mask $\mathbf{M}_{p}$ to identify the source bag for each instance within pack $p$. The embedding for the $b$-th bag is then computed by selectively aggregating its instances from the pack: $z^{\mathrm{main}}_{b}=\Gamma_{\theta}\!\Bigl(\mathbf{M}_{pb},P^{\mathrm{main}}_{p}\Bigr)$. We compute the main loss $\mathcal{L}_{\text{main}}$ over $B$ slides, based on slide-level predictions $\hat{y}_{b}=g_{\phi}(z^{\mathrm{main}}_{b})$. For the residual branch, each pack $p$ acts as a hyperslide, introducing high-level supervision. We compute its embedding $z^{\mathrm{res}}_{p}=\Gamma_{\theta}\!\Bigl(P^{\mathrm{res}}_{p}\Bigr)$, obtain pack-level predictions $\hat{y}^{\mathrm{hyper}}_{p}=g_{\phi}(z^{\mathrm{res}}_{p})$, and compute the residual loss $\mathcal{L}_{\text{res}}$ over $B^{\prime\prime}$ packs. The overall objective function is a weighted sum of the two losses: $\mathcal{L}=\mathcal{L}_{\text{main}}+\lambda\,\mathcal{L}_{\text{res}}$.
@@
Packing. The packing operation processes concatenated downsampled embeddings, denoted as $\mathcal{H}=\bigcup_{b=1}^{B}\tilde{\mathcal{R}}_{b}=\{h_{n}\}_{n=1}^{M}$, where $M$ is total number of features in the mini-batch and $\beta_{n}$ is original bag index of feature $h_{n}$. The operation sequentially fills packs of fixed length $L$. Let $P_{p}\in\mathbb{R}^{L\times D}$ be the $p$-th pack, for $p=1,\dots,B^{\prime}$. Features $\{h_{n}\}_{n=1}^{M}$ are placed sequentially into $P_{1},P_{2},\dots,P_{B^{\prime}}$. When a pack $P_{p}$ cannot accommodate the next feature without exceeding $L$ tokens, or when all $M$ features have been placed, any remaining positions in $P_{p}$ are filled with vectors $\mathbf{0}\in\mathbb{R}^{D}$.
@@
Isolated Mask. We employ auxiliary masks to process features within each pack $P_{p}$ while preserving original bag integrity. Based on the CPath pipeline above, masks are divided into aggregation-oriented and classification-oriented. Aggregation masks constrain the feature aggregation stages, ensuring that computations within each pack only involve instances from the corresponding source bag. For example, in Transformer-based models like TransMIL Shao et al. (2021), this mask restricts Multi-Head Self-Attention computations strictly to features originating from the same source bag within the pack, preventing cross-bag attention and aggregation. Classification-oriented masks subsequently select tokens relevant for prediction. This mask identifies source bag for each non-padding token, enabling appropriate classification based on bag identity. These tailored masks enable efficient and effective MIL within our pack-based framework. Detailed definition is provided in Appendix D.2.
@@
Attention-driven Downsampler (ADS). WSIs exhibit significant redundancy  Tang et al. (2023), Zhang et al. (2024b), which can also manifest between the feature sets derived from $\mathcal{R}_{b}$ and $\mathcal{D}_{b}$. To fuse these features while mitigating redundancy in dual branches, we designed the ADS module. This module utilizes attention-driven downsampling to produce a compact and informative representation. Given a set of $N$ embeddings $\{h_{i}\}_{i=1}^{N}$, ADS first computes a per‐instance attention score via a shallow MLP and applies it as a residual weigh $u_{i}=h_{i}+a_{i}\,h_{i}$. Followed by a learnable linear layer, we compute $W^{L}\!\in\!\mathbb{R}^{D\times D}$, $v_{i}=u_{i}\,W^{L}$. To facilitate structured downsampling that preserves spatial relationships, we perform instance unshuffle, which rearranges the sequential features into a group-based representation:
@@
where $k$ is the downsampling factor. ADS poolings along the $k$ dimension of each pack, followed by a projection $W^{P}\!\in\!\mathbb{R}^{D\times D}$. ADS takes the input features $\{h_{i}\}$ and a factor $k$, resulting in the output sequence $\operatorname{ADS}(\{h_{i}\};k)=\{\tilde{h}_{j}\}_{j=1}^{\lceil N/k\rceil}\in\mathbb{R}^{\lceil N/k\rceil\times D}$. Each downsampled feature $\tilde{h}_{j}$ is computed as:
@@
where $\operatorname{Pool}(\cdot)$ is either random or max pooling, and $W^{P}\!\in\!\mathbb{R}^{D\times D}$ is a projection matrix. By weighting instances with attention scores $a_{i}$} and pooling across structured groups, ADS reduces the instance count by a factor of $k$ while prioritizing regions of high clinical relevance. ADS also maintains interpretability at inference time. Additional details and a discussion on the applicable boundaries of ADS are provided in Appendix B.3.
Inference Pipeline. For inference, we adopt a deterministic pipeline to ensure reproducible predictions. The residual branch and stochastic sampling are bypassed. Each slide is processed individually with batchsize of 1, feeding its complete sequence of instance embeddings into the main branch.
Figure 4: Illustration of Task-specific Hyperslide Labels.

### v2

# Revisiting Data Challenges of Computational Pathology: A Pack-based Multiple Instance Learning Training Framework

@@ Abstract
Computational pathology (CPath) digitizes pathology slides into whole slide images (WSIs), enabling analysis for critical healthcare tasks such as cancer diagnosis and prognosis. However, WSIs possess extremely long sequence lengths, significant length variations (from 200 to 200K), and limited supervision. These extreme length variations lead to high data heterogeneity and redundancy. Conventional methods often compromise on training efficiency and optimization to preserve such heterogeneity under limited supervision. To comprehensively address these challenges, we propose a pack-based MIL framework. It packs multiple sampled, variable-length feature sequences into fixed-length ones, enabling batched training while preserving data heterogeneity. Moreover, we introduce a residual branch that composes discarded features from multiple slides into a hyperslide which is trained with tailored labels. It offers multi-slide supervision while mitigating feature loss from sampling. Meanwhile, an attention-driven downsampler is introduced to compress features in both branches to reduce redundancy. By alleviating these challenges, our approach achieves an accuracy improvement of up to 8% while using only 12% of the training time in the PANDA (UNI). Extensive experiments demonstrate that focusing data challenges in CPath holds significant potential in the era of foundation models. The code is here.
@@ 1 Introduction
Computational pathology (CPath) [song2023artificial, cifci2023ai] represents a rapidly evolving interdisciplinary research domain that integrates advanced computer vision techniques and pathology to facilitate accurate and efficient interpretation of histopathological images. Central to CPath are whole slide images (WSIs, slides), digitized pathology slides with gigapixel resolution. It enables comprehensive microscopic analysis to support critical healthcare tasks such as cancer sub-typing [ilse2018attention, zhang2022dtfd, tu2022dual], grading [panda], and prognosis [wen2023deep, yao2020whole]. Patching strategies help researchers effectively process these gigapixel images within hardware constraints. However, patches derived from WSIs present following challenges: as shown in Fig.1, 1) they possess extremely long sequence lengths and significant sequence length variations (e.g., from 200 to 200K in TCGA-BRCA-Survival). Such extreme distributions in sequence length, coupled with diverse morphological characteristics, contribute to data heterogeneity, which is substantial for CPath tasks. 2) and introduce input redundancy challenges for CPath algorithms. 3) Moreover, due to the gigapixel resolution and specialization, WSIs typically have only slide-level annotations, lacking more supervision that matches the complex input.
@@
Current two-stage multiple instance learning (MIL) [maron1997mil_1] paradigm [clam] is a compromise resulting from high data heterogeneity and limited supervision. This paradigm employs a pre-trained encoder to extract offline patch (instance) features, and uses a MIL model to produce slide-level (bag-level) results. Due to data challenges, it suffers from training inefficiency and instability. Specifically, with significant variations in sequence length across slides, mainstream methods typically process data with a batchsize of 1 during training [shao2021transmil, li2024dynamic]. While these approaches preserve whole-slide heterogeneity, training with a batchsize of 1 is inefficient (e.g., training TransMIL [shao2021transmil] on the PANDA [panda] dataset requires over 50 RTX3090 GPU-hours) and may yield suboptimal performance [koga2025attention]. A few methods [campanella2019clinical, liu2024attention] attempt to enable batched training by sampling or padding all sequences to a uniform length; however, this approach can lead to a loss of data heterogeneity and important features, especially affecting complex methods and tasks.
@@
To comprehensively address three data challenges, we propose a novel pack-based MIL framework. Inspired by recent advancements in large model [pouransari2024dataset, krell2021efficient, dehghani2023patch, wang2024qwen2], it packs multiple variable-length sequences into a single fixed-length sequence to enable batched training while preserving data heterogeneity. However, leveraging packing strategies for effective batched training in CPath is far from straightforward. The excessive length of packed sequences hinders training, necessitating patch sampling, which still leads to feature loss. Therefore, we split the input features into main and residual branches, packing the kept and discarded features, respectively, to minimize sampling-induced feature loss. In the main branch, we employ masks to maintain the independence of different slides within a pack. Conversely, the residual branch treats discarded features from multiple slides in the same pack as a single hyperslide. To train this hyperslide effectively, we introduce task-specific hyperslide labels and loss functions. Crucially, this approach effectively offers multi-slide supervision.
@@
While some outstanding works have explored supplementary supervision [zhang2022dtfd, shao2023lnpl, brussee2025graph, fang2024sam], most focus on intra-slide modeling (e.g., instance-level or pseudo-bag), neglecting inter-slide relationships. Pathology slides from the same spatial and tissue origin exhibit consistent morphological characteristics [lin2025impact, chen2022fast, kaczmarzyk2024open]. Learning inter-slide correlations allows the hyperslide to provide the model with a more comprehensive perspective, enabling the discovery of more generalizable pathological features. Furthermore, we propose an attention-driven downsampler to compress features for reducing input redundancy within both branches. To validate our framework, we conducted extensive experiments using features from foundation models. Results demonstrate that our approach consistently improves multiple baselines by effectively mitigating the data challenges inherent in CPath. Specifically, it delivers substantial performance gains (e.g., +11% accuracy on PANDA) while improving training efficiency ($\sim 8\times$ speedup on PANDA). Our contributions are:
@@ 2 Related Works
Supervision in Computational Pathology. Recent CPath advancements leverage MIL to reduce annotation burden. Using only slide labels, MIL has evolved with mechanisms like attention [ilse2018attention, tang2023mhim, zhang2024attention, zhang2025aem, dong2025fast], clustering [lin2023interventional], Transformers [shao2021transmil, jaume2024transcriptomics, fourkioti2023camil], and GNNs [wang2021hierarchical, eastwood2023mesograph] to improve interpretability and accuracy. Complementing pure MIL methods, pseudo-labeling strategies have emerged as powerful techniques, encompassing instance-level pseudo-labeling [qu2022dgmil], knowledge distillation frameworks [zhang2022dtfd, qu2022bi], limited pathologist patch annotations [koga2025attention], weak regional annotations [wang2022label], and semi-supervised consistency regularization [jiang2023semi]. These hybrid strategies effectively generate additional supervision to refine instance predictions and leverage unlabeled data, boosting performance and data efficiency. While some studies [liu2024pseudo, ouyang2024mergeup, aswolinskiy2025attention] explore mixup-like data augmentation between WSI pairs, supervision leveraging relationships across multiple slides is still unexplored.
Batchsize in Computational Pathology. Batchsize is a crucial hyperparameter in deep learning. However, its exploration in CPath remains limited, primarily due to the computational demands of WSIs and the inherent data heterogeneity within each slide. Consequently, mainstream slide-level MIL methods typically adopt a batchsize of 1 [li2024generalizable, shi2024vila, li2024dynamic, song2024morphological, zhang2024attention, li2024rethinking, fourkioti2023camil, tang2025multipleinstancelearningframework, tang2025revisitingendtoendlearningslidelevel]. For instance, RRTMIL [tang2024feature] utilizes slide-wise regional and cross-region self-attention to capture patch ordinality and heterogeneity within each slide, necessitating a batchsize of 1 to maintain intra-slide relationships. Despite its prevalence, this practice often results in training instability and slow convergence, prompting methods such as gradient accumulation over multiple slides [koga2025attention, zhang2025icfnet] or instance-level sampling strategies that select fixed-size subsets of patches per slide to mitigate computational overhead and improve learning stability [campanella2019clinical, liu2024attention]. Current slide-level methods treat batched inputs as an efficiency trade-off rather than a genuinely effective training strategy. In this paper, we investigate the benefits of pack-based batched training for slide‑level prediction and explore practical guidelines in CPath. Supplementary gives more discussion about this topic and pack-based training.
@@ 3 Method
3.1 MIL-based Computational Pathology
@@
Histopathological WSIs are often gigapixel resolution, making direct processing impractical. Current approaches typically use weakly supervised MIL, where a WSI is treated as a bag $\mathcal{X}=\{x_{1},\dots,x_{N}\}$ of instances. During training, only a slide‑level label $y$ is available. Each instance $x_{i}$ is encoded to an embedding $h_{i}=f(x_{i})$ using a offline feature extractor $f$. An aggregation function $\Gamma(\cdot)$ combines instance embeddings $\{h_{i}\}_{i=1}^{N}$ into a bag-level representation $z=\Gamma_{\theta}(\{h_{i}\}_{i=1}^{N})$. This representation $z$ is then used by a classifier $g_{\phi}$ to predict the slide label $p(y\,|\,\mathcal{X})=g_{\phi}(z)$. The significant variation in instance count $N$ across WSIs contribute to data heterogeneity. This heterogeneity along with its associated spatial and morphological context, is crucial for CPath, as shown in Fig.2. Specifically, when all instances are randomly sampled to a fixed-length sequence, a latest method like RRTMIL [tang2024feature] exhibits consistent performance degradation on multiple benchmarks, particularly for complex tasks like survival analysis. To maintain data heterogeneity, current methods necessitates training with a batchsize of 1, resulting in noisy gradient estimates and optimization instability.
Figure 2: Impact of data heterogeneity on CPath. RS represents random sampling instances in all bags to a fixed length while maintaining original label, thus losing data heterogeneity.
Figure 3: Left: Overview of proposed pack-based MIL training framework. Instance features from each WSI are sampled into kept and discarded sequences. Both sequences are processed by ADS. Downsampled sequences from different bags are then concatenated into fixed-length packs. This packing mechanism aggregates different WSIs into fixed-length packs, thereby enabling batched training. The dual-branch architecture, with shared weights, processes: 1) The Main Branch, supervised by slide-level labels, 2) The Residual Branch, supervised by Hyperslide labels. Right: Architecture of the Attention-driven Downsampler (ADS). Pseudo-codes are in Supplementary.
3.2 Pack-based MIL Training Framework
@@
CPath poses significant data challenges, including high data heterogeneity, redundancy and limited supervision, which hinder mainstream MIL method. To overcome these issues, we propose a pack-based MIL training framework, named PackMIL. As illustrated in Fig.3, the proposed method effectively employs a packing operation to maintain data heterogeneity and allow batched input, and it further provides multi-slide supervision, leading to more effective and stable training. Consider a mini‑batch of $B$ bags, the bb‑th bag provides a set of instance embeddings $\mathcal{H}_{b}=\{h_{bi}\}_{i=1}^{N_{b}}$ where $h_{bi}\!\in\!\mathbb{R}^{D}$. However, the feature sequences extracted from gigapixel WSIs can be prohibitively long, making their direct inclusion into fixed-length packs impractical due to memory and computational limitations during training. To address this, we employ stochastic instance-level sampling with ratio $r\in(0,1)$. The subsets $\mathcal{K}_{b}$ and $\mathcal{D}_{b}$ are formed as:
${\mathcal{K}}_{b}=\{\,h_{bi}\mid m_{bi}=1\,\},\quad\mathcal{D}_{b}=\{\,h_{bi}\mid m_{bi}=0\,\},$ (1)
@@
where $m_{bi}\sim\operatorname{Bernoulli}(1-r)$. We introduce the Residual Branch to mitigate feature loss during sampling and establish hyperslide supervision. The kept sets $\mathcal{K}_{b}$ (main branch) and discarded sets $\mathcal{D}_{b}$ (residual branch) are processed to produce downsampled feature sets, $\tilde{\mathcal{K}}_{b}$ and $\tilde{\mathcal{D}}_{b}$, respectively. To maintain data heterogeneity for batch training, features from sampled sets of each bag are sequentially arranged into fixed-length packs of size $L$. This packing operation $\operatorname{PACK}(\cdot)$ processes all bags in the mini-batch to form:
@@
$\begin{split}\mathcal{P}^{\mathrm{main}}&=\operatorname{PACK}\Bigl(\{\tilde{\mathcal{K}}_{b}\}_{b=1}^{B},\,L\Bigr)\in\mathbb{R}^{B^{\prime}\times L\times D},\\ \mathcal{P}^{\mathrm{res}}&=\operatorname{PACK}\Bigl(\{\tilde{\mathcal{D}}_{b}\}_{b=1}^{B},\,L\Bigr)\in\mathbb{R}^{B^{\prime\prime}\times L\times D},\end{split}$ (2)
@@
We also incorporate an attention-driven downsampler (ADS) module for handling input redundancy between the two branches. This module fuses features from instances in $\mathcal{K}_{b}$ and $\mathcal{D}_{b}$ to generate more compact and informative feature representations. Each set $\mathcal{S}\in\{\mathcal{K}_{b},\mathcal{D}_{b}\}$ is downsampled by a factor $k\in\mathbb{N}$ using an ADS module, yielding $\tilde{\mathcal{S}}=\operatorname{ADS}(\mathcal{S};k)$ with size $|\tilde{\mathcal{S}}|=\lceil|\mathcal{S}|/k\rceil$.
@@
After packing, the main and residual branches are processed independently while sharing the same MIL model weight. In the main branch, each pack contains features originating from $B$ distinct bags. We use a mask $\mathbf{M}_{p}$ to identify the source bag for each instance feature within $p$-th pack. The embedding for the $b$-th bag is then computed by aggregating its instance features from the pack: $z^{\mathrm{main}}_{b}=\Gamma_{\theta}\!\Bigl(\mathbf{M}_{pb},\mathcal{P}^{\mathrm{main}}_{p}\Bigr)$. We compute the main loss $\mathcal{L}_{\text{main}}$ over $B$ slides, based on slide-level predictions $\hat{y}_{b}=g_{\phi}(z^{\mathrm{main}}_{b})$. For the residual branch, each pack $p$ acts as a hyperslide, introducing high-level supervision. We compute its embedding $z^{\mathrm{res}}_{p}=\Gamma_{\theta}\!\Bigl(\mathcal{P}^{\mathrm{res}}_{p}\Bigr)$, obtain pack-level predictions $\hat{y}^{\mathrm{hyper}}_{p}=g_{\phi}(z^{\mathrm{res}}_{p})$, and compute the residual loss $\mathcal{L}_{\text{res}}$ over $B^{\prime\prime}$ packs. The overall objective function is a weighted sum of the two losses: $\mathcal{L}=\mathcal{L}_{\text{main}}+\lambda\,\mathcal{L}_{\text{res}}$.
@@
Packing. The packing operation processes concatenated downsampled embeddings, $\tilde{\mathcal{K}}_{b}$ and $\tilde{\mathcal{D}}_{b}$. The operation sequentially fills packs of fixed length $L$. Let $\mathcal{P}_{p}\in\mathbb{R}^{L\times D}$ be the $p$-th pack, for $p=1,\dots,B^{\prime}$. Embeddings are placed sequentially into $\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{B^{\prime}}$. When a pack $\mathcal{P}_{p}$ cannot accommodate the next feature without exceeding $L$, or when all feature have been placed, any remaining positions in $\mathcal{P}_{p}$ are filled with vectors $\mathbf{0}\in\mathbb{R}^{D}$. To enforce the fixed pack length $L$, any downsampled features still exceeding this length are truncated to $L$. Packing pseudocode is provided in Supplementary.
@@
Isolated Mask. We employ auxiliary masks to process features within each pack $\mathcal{P}_{p}$. These masks serve a dual purpose: 1) preserving bag integrity by preventing cross-bag interactions, and 2) nullifying the impact of zero-padding. Based on the CPath pipeline above, masks are divided into aggregation-oriented and classification-oriented. Aggregation masks constrain the feature aggregation stages, ensuring that computations within each pack only involve instance features from the same source bag. Classification-oriented masks subsequently select non-padding features for the final prediction. These tailored masks enable efficient and effective MIL within our pack-based framework. Detailed definition is provided in Supplementary.
Figure 4: Illustration of Task-specific Hyperslide Labels. A hyperslide is a conceptual pack formed by aggregating multiple WSIs to enable inter-slide supervision. These task-specific labels are constructed based on the clinical characteristics of different clinical tasks.
@@
Attention-driven Downsampler (ADS). WSIs exhibit significant redundancy  [tang2023mhim, zhang2024attention], which can also manifest between the feature sets derived from $\mathcal{K}_{b}$ and $\mathcal{D}_{b}$. To fuse features while mitigating redundancy in dual branches, we designed the ADS module. This module utilizes attention-driven downsampling to produce a compact and informative representation. Given a set of $N$ embeddings $\{h_{i}\}_{i=1}^{N}$, ADS computes a per-instance normalized attention score via a shallow MLP and applies it as a residual weigh $u_{i}=h_{i}+a_{i}\,h_{i}$, which preserves the original feature while enhancing it based on its importance. Followed by a learnable linear layer, we compute $W^{L}\!\in\!\mathbb{R}^{D\times D}$, $v_{i}=u_{i}\,W^{L}$. Then , we perform instance unshuffle, which rearranges sequential instance features into group-based representation:
@@
where $k$ is the downsampling factor. ADS performs pooling along the $k$ dimension of each pack, followed by a projection $W^{P}\!\in\!\mathbb{R}^{D\times D}$. ADS takes the input features $\{h_{i}\}$ and a factor $k$, resulting in the output sequence $\operatorname{ADS}(\{h_{i}\};k)=\{\tilde{h}_{j}\}_{j=1}^{\lceil N/k\rceil}\in\mathbb{R}^{\lceil N/k\rceil\times D}$. Each feature $\tilde{h}_{j}$ is computed as:
@@
where $\operatorname{Pool}(\cdot)$ is either random or max pooling, and $W^{P}\!\in\!\mathbb{R}^{D\times D}$ is a projection matrix. By weighting instances with attention scores $a_{i}$} and pooling across structured groups, ADS reduces the instance count while prioritizing regions of high clinical relevance , as the attention is trained with task supervision. ADS also maintains interpretability at inference time. Additional details, discussions on applicable boundaries, and pseudocode are in the Supplementary.
Inference Pipeline. For inference, we adopt a deterministic pipeline to ensure reproducibility. The residual branch and stochastic sampling are bypassed. Each slide is processed individually with batchsize of 1, feeding its complete sequence of instance embeddings into the main branch.